OpenStudy (anonymous):

Solve. 7x − 2(3x + 4) =15?? show work

OpenStudy (anonymous):

first distribute the 2 into the parenthess

OpenStudy (anonymous):

you should get: 7x-6x-8=15

OpenStudy (anonymous):

now, collect like terms

OpenStudy (saifoo.khan):

7x -6x-8 = 15 x -8 = 15 x = 15 + 8 x = 23

OpenStudy (anonymous):

x-8=15 then add 8 to both sides x=23

OpenStudy (anonymous):

you can reward people that answer by clicking the good answer button